Determine whether each of these functions is a bijection from R to R?f (x) = (x2 + 1)/(x2 + 2)


1
Expert's answer
2021-09-07T19:04:09-0400

Let us determine whether the function "f:\R \to \R,\ f (x) = \frac{x^2 + 1}{x^2 + 2}," is a bijection. Taking into account that for "x_1=-1" and "x_2=1\ne x_1" we get

"f(x_1)=f(-1)= \frac{(-1)^2 + 1}{(-1)^2 + 2}=\frac{2}{3}=\frac{1^2 + 1}{1^2 + 2}=f(1)=f(x_2),"

we conclude that the function "f" is not an injection, and hence this function is not a bijection.